OpenStudy (anonymous):

Physics help! A BB-gun is aimed directly at the center of a target 50m away. If the BB is fired horizontally at 350m/s how far below the bulls-eye wil the BB hit?

OpenStudy (anonymous):

are you familiar with kinematic equations

OpenStudy (anonymous):

first find the time that the bullet takes to get to the target, this is in the x direction we know V = d / t so, t =d / v once we have time, we can put it into a kinematic equation to find the distance the bullet wil fall use the equation d = 1/2*a*t^2 where d is the answer to the question, a is acceleratin due to gravity, and t is the time caculated in the first part.

OpenStudy (anonymous):

Okay, I got .142 seconds for the time

OpenStudy (anonymous):

but I'm not sure if that's right because I plugged it in with that second equation and i got .0988036

OpenStudy (anonymous):

ok ill work it out and give you my answer

OpenStudy (anonymous):

I got pretty much the same, 0.1m, which is reasonable distance for the bullet to drop

OpenStudy (anonymous):

10cm seems about right

OpenStudy (anonymous):

thank you so much for your help!